SAT I scores around the nation tend to have a mean scale score around 500, a standard deviation of about 100 points, and are approximately normally distributed. What SAT I score within the population would have a percentile rank of approximately 2.5? Show all work as to how this is obtained.

Answer by stanbon(75887)

SAT I scores around the nation tend to have a mean scale score around 500, a standard deviation of about 100 points, and are approximately normally distributed. What SAT I score within the population would have a percentile rank of approximately 2.5? Show all work as to how this is obtained.
---
Find the z-value with a left tail of 0.025
invNorm(0.025) = -1.96
----
Find the score using x = z*s + u
x = -1.96*100+500 = 500-196 = 304
====================
Cheers,
Stan H.